OpenStudy (anonymous):

IS the answer 0 here

OpenStudy (anonymous):

If log7log2logpi(x) vanishes , then x equals (

ganeshie8 (ganeshie8):

that equation is begging for using latex^

OpenStudy (anonymous):

Yes wait my computer is misbehaaving

OpenStudy (anonymous):

\(\large \log7\log2\log \pi(x)\)? or \(\large \log7(\log2(\log \pi(x)))\)?

OpenStudy (shadowlegendx):

or Bacon :O

ganeshie8 (ganeshie8):

@BSwan use `a_0` for getting \(a_0\)

OpenStudy (shadowlegendx):

\(\bbox[5pt, #33fd33 ,border:5px solid #872af0]{\bf\huge\ It\ Is\ Bacon!!!}\)

OpenStudy (shadowlegendx):

;-;

OpenStudy (anonymous):

I will post a new post . My computer is not functioning properly

OpenStudy (shadowlegendx):

Ok, we shall assist you there